Sam sells industrial equipment. He converts one in four customers he cold calls seeking an appointment. He closes on one out of every ten customers he visits. His goal is to earn $96,000 per year, allocated equally across each month. His commission rate is $800 for each item sold.

Allowing for 20 working days per month, on average how many daily sales calls each month must Sam make to earn $96,000?

a) 20
b) 30
c) 40
d) 120
e) None of the above
Answer #1
Appointment to sales call ratio = 1/4
Sales to appointment ratio = 1/10
Number of items to be sold each year to earn 96000 = 96000 / Commission rate per item = 96000 / 800 120
Total number of sales call needed per year to sell 120 items = 120 / Sales to appointment ratio / Appointment to sales call ratio = 120 / (1/10) / (1/4) 4800
Daily sales call needed = Total number of sales call needed per year to sell 120 items / ( 12 * Working days in a month ) = 4800 / ( 12 * 20 ) 20 option a